### v2

@@ 1 Introduction
The paper is organized as follows. In Section 2, we recall the geometric setting for photon surfaces in spherically symmetric spacetimes and derive their evolution equations. Section 3 introduces the marginally bound LTB collapse model and discusses how the photon surface condition adapts to this context. We analyze the possible extensions of the photon surface from the Schwarzschild exterior into the collapsing interior, distinguishing between spacelike extensions and those corresponding to null radial geodesics. In Section 4, we focus on the qualitative behavior of these surfaces, examining their ability to shield the central singularity. We show that the nature of the singularity—naked or covered—dictates whether the photon surface terminates at the regular center or reaches the singularity itself. Finally, Section 5 summarizes the main findings and suggests possible directions for further investigation.
@@ 2 Photon surfaces in spherical symmetry
where $\lambda$, $\nu$, and $r$ are functions of $(t,x)$ and $\mathrm{d}\Omega^{2}=\mathrm{d}\theta^{2}+\sin^{2}\theta\,\mathrm{d}\phi^{2}$. We shall proceed to apply Theorem 1. Since $S$ is timelike, we can parameterize it with coordinates $(t,x(t),\theta,\phi)$, with $x(t)$ to be determined from the equation (1). In this situation, indeed, the area of $\mathcal{T}_{s}$ is given by $A_{s}=4\pi x(t(s))$, and (1) takes the form
@@
where $\dot{x}(t)$ and $\ddot{x}(t)$ stand for derivatives of $x$ with respect to $t$ and, given a function $f(t,x)$, $f_{t}$ and $f_{x}$ denote its partial derivatives. We can transform this second-order ODE into a first-order system of ODEs as follows
@@
Let us consider the particular case where (2) is static, therefore admitting a $SO(3)\times\mathbb{R}$ group of isometries, where the $\mathbb{R}$ isometries are generated by a Killing vector field $K$, orthogonal to the $SO(3)$ orbits. In this context, the notion of photon sphere is defined in [Claudel:2000yi] as an $SO(3)\times\mathbb{R}$ invariant photon surface. In the static case, $\nu,\lambda$ and $r$ are functions of $x$ only, and the Killing vector field that generates the isometries of $\mathbb{R}$ is collinear to $\partial_{t}$. Therefore, a photon space is a photon surface with constant $x(t)$, which in view of (4a)–(4b) is implicitly defined by those $x$ satisfying
@@
3 Photon surfaces in dust collapse
Now, let us apply the notions sketched in the previous section to the well-known model of a spherical dust cloud undergoing gravitational collapse. This is obtained by considering a general spherical metric (2) where the energy–momentum tensor is given by $T=-\rho\,\mathrm{d}t\otimes\partial_{t}$ [Singh:1994tb, Joshi:2008zz]. The metric is given by the Lemaître–Tolman–Bondi (LTB) model
@@ 3 Photon surfaces in dust collapse
The initial data $k(x)$ and $m(x)$ are prescribed at some initial time $t_{0}$, i.e., $t_{0}=0$, together with the initial energy $\rho_{0}(x):=\rho(t_{0},x)$. The value of the Misner-Sharp mass $m(x)=\frac{r}{2}\left(1-g^{\alpha\beta}r_{\alpha}r_{\beta}\right)$ will also determine the initial condition $r(t_{0},x)$, because integrating the Einstein field equation $m_{x}=4\pi\rho r^{2}r_{x}$ along $t=t_{0}$ one obtains
@@
As pointed out in [Cao:2019vlu, (5.5)], for the LTB model a lengthy but straightforward computation transforms the equation (3) for photon surfaces as
which can be given in the form of the 1st-order system (4a)–(4b):
$\dot{x}(t)$ $=\frac{\sqrt{k(x)+1}}{r_{x}}\,y(t),$ (10a)
When it comes to assigning initial conditions for this system, we first have to observe that the metric (7) describes only the interior region of the model, i.e., the dust star. In other words, this metric is defined for $x\in[0,x_{b}]$, where $x_{b}$ represents the boundary of the star. The timelike hypersurface $\Sigma=\{x=x_{b}\}$ must be matched with a suitable exterior solution that describes the matter distribution outside the spherical dust cloud. It is well known that for the exterior region to be Schwarzschild spacetime, the radial pressure, indicated by the $T^{1}_{1}$ component of the energy-momentum tensor, must be zero at $\Sigma$. According to the Einstein field equations, this condition ensures that the Misner–Sharp mass does not depend on time $t$ and remains constant across the matching hypersurface, a necessary condition for matching [Magli_1997]. Naturally, this applies to the model being analyzed here, where pressures uniformly vanish throughout the spacetime.
@@
We will refer to this model as the marginally bound LTB collapse. Its end state, starting from regular initial conditions at $t=0$, is well known in the literature [Singh:1994tb]: let us summarize it briefly. Consider a mass profile such that, using (11)
@@
$\{(t,x)\,:\,x\in[0,x_{b}],\,\,t<t_{s}(x)\},$
matched to a Schwarzschild exterior
at the star boundary $\Sigma=\{x=x_{b}\}$. Israel–Darmois junction conditions [Israel:1966rt] is sufficient to ensure that the matching is $C^{1}$ across $\Sigma$, provided that the Schwarzschild mass is given by $M=m(x_{b})$. Observe that the coordinates $(\tau,r)$ in (13) will be related to $t$ on the matching surface – in particular, $r=r(t,x_{b})$ on $\Sigma$ [^2] That justifies the slight abuse of notation where we employ the same letter $r$ for both the radial coordinate in (13) and the metric coefficient (11) within (7).. Details can be found in [Giambo2005GravitationalFields].
Let us now consider the photon surface for this model. On the exterior (13), of course, the photon surface is given by $r=3M$. When it comes to determining its extension to the interior spacetime, we need to set initial conditions for the system (10a)–(10b). The first condition is necessarily given by $x(t_{0})=x_{b}$, where $t_{0}$ satisfies the equation $r(t_{0},x_{b})=3m(x_{b})$, i.e., using (11),
@@
Of course, one may consider other extensions that start spacelike at $x=x_{b}$. These choices correspond to setting $|y_{0}|<1$. However, $y_{0}<0$ would correspond to a decreasing $x(t)$ curve and would, for the same reason, be less significant than the $y_{0}=1$ extension. The following lemma will help us rule out all the other cases.
@@
The marginally bound LTB collapse admits an extension of the photon surface $r=3M$ in the interior of the cloud. This is the only physically acceptable extension, and it is given by the surface
@@
4 Photon surface analysis
@@
The curve $t_{\xi}(x)$ is a supersolution of (18) and $x_{b}$ chosen such that $t_{0}\in]t_{\xi}(x_{b}),t_{h}(x_{b})[$. The solution of (18) can be extended back for $x<x_{b}$: it cannot cross the subsolution $t_{h}(x)$ from below nor the supersolution $t_{\xi}(x)$ from above and then it must extend up to the central singularity $(x=0,t=1)$.
@@
In the marginally bound LTB collapse, the photon surface extends back to the central singularity if and only if it is naked. In this case, there exist null radial geodesics $t_{g}(x)$ emanating from the central naked singularity such that $t_{g}(x)<t(x)$, $\forall x\in]0,x_{b}]$, where $t(x)$ is the solution of (18) associated with the photon surface extension established in Theorem 2.
@@ 5 Discussion and conclusions
Our results indicate that the only physically consistent extension of the photon surface $r=3M$ into the collapsing cloud is given by a surface generated by outgoing null radial geodesics. This extension satisfies the photon surface equation and adheres to the junction conditions at the boundary with the Schwarzschild exterior. Importantly, the nature of the central singularity—whether it is covered or naked—plays a crucial role in determining the global structure of the photon surface. When the singularity is covered, the photon surface terminates at the regular center before the singularity forms. Conversely, if the singularity is naked, the photon surface can extend all the way back to the central singularity itself. However, in such cases, the surface is not sufficient to prevent the escape of photons from the singularity, highlighting its limited covering capability. Previous studies have established that naked singularities create different lensing effects compared to black holes [Virbhadra:2022iiy]. Additionally, research by Ortiz et al. [Ortiz:2015rma] demonstrates differences in the shadow evolution between naked singularities and black holes. While both objects eventually appear similar to distant observers, their formation dynamics differ significantly, also due to the infinitely redshifted photons that escape from the central naked singularity to the external region – a phenomenon documented in earlier work [Dwivedi:1998ts, Giambo:2006vr]. Our findings in the present paper support these observations by showing that shadow evolution also varies in terms of photon surface extension, suggesting that understanding photon surface geometry is crucial for distinguishing these compact objects, as the formation and features of the shadow are determined by the whole underlying spacetime structure.
@@ Acknowledgement.
The authors thank LM Cao and Y Song for their insightful discussions.

### v3

@@ 1 Introduction
Before proceeding, it is worth stressing that horizons and photon surfaces, although both tied to the causal structure of spacetime, address different physical questions. In the context of gravitational collapse, an apparent horizon is the outermost marginally trapped 2-sphere: the surface on which outgoing null geodesics have vanishing expansion, so that light directed outward neither spreads nor converges [PhysRevLett.14.57, Hawking:1973uf]. It marks the boundary of the trapped region, beyond which even outward-pointing photons are dragged inward. A photon surface, by contrast, is a hypersurface with the property that every null geodesic initially tangent to it remains tangent under propagation [Claudel:2000yi]; it captures the locus where photons can stay confined to the surface. In Schwarzschild spacetime the photon sphere envelops the event horizon, located in a stronger gravitational regime. We shall see that the same hierarchy persists in a dynamical setting also, and we will highlight how the two surfaces evolve differently during collapse.
The paper is organized as follows. In Section 2, we recall the geometric setting for photon surfaces in spherically symmetric spacetimes and derive their evolution equations. Section 3 provides an overview of the fundamental concepts related to spherical dust collapse and reviews the key features of the singularity emerging at the end of the collapse. Section 4 focuses on determining the photon surface for the model at hand and showing that the nature of the singularity—naked or covered—dictates whether the photon surface terminates at the regular center or reaches the singularity itself. Finally, Section 5 summarizes the main findings and suggests possible directions for further investigation.
@@ 2 Photon surfaces in spherical symmetry
where $\lambda$, $\nu$, and $r$ are functions of $(t,x)$ and $\mathrm{d}\Omega^{2}=\mathrm{d}\theta^{2}+\sin^{2}\theta\,\mathrm{d}\phi^{2}$. We shall proceed to apply Theorem 1. Since $S$ is timelike, we can parameterize it with coordinates $(t,x(t),\theta,\phi)$, with $x(t)$ to be determined from the equation (1). In this situation, indeed, the area of $\mathcal{T}_{s}$ is given by $A_{s}=4\pi\mathbf{r(t,x(t))^{2}}$, and (1) takes the form
@@
where $\dot{x}(t)$ and $\ddot{x}(t)$ stand for derivatives of $x$ with respect to $t$ and, given a function $f(t,x)$, $f_{t}$ and $f_{x}$ denote its partial derivatives.
For the purpose of the present paper, the second-order ODE (3) is not the best way to study the photon surface and our strategy here is to shift the perspective from a purely geometric PDE to a dynamical systems framework: this crucial step is formalized in the following Proposition.
Proposition 2.
Let $S$ be a photon surface for the metric (2) parameterized by $(t,x(t),\theta,\phi)$. Then, the second-order PDE (3) is strictly equivalent to the following first-order non-autonomous dynamical system:
@@
We stress that the reformulation provided by the above Proposition is not merely a formal manipulation. By recasting (3) into the first-order system (4a)–(4b), we reveal the underlying non-autonomous dynamical nature of the problem. What’s more, this structure is decisive to prove the unique extension of the photon surface into the collapsing dust cloud, as we will see in Section 4.
@@
Let us consider the particular case where (2) is static, therefore admitting a $SO(3)\times\mathbb{R}$ group of isometries, where the $\mathbb{R}$ isometries are generated by a Killing vector field $K$, orthogonal to the $SO(3)$ orbits. In this context, the notion of photon sphere is defined in [Claudel:2000yi] as an $SO(3)\times\mathbb{R}$ invariant photon surface. In the static case, $\nu,\lambda$ and $r$ are functions of $x$ only, and the Killing vector field that generates the isometries of $\mathbb{R}$ is collinear to $\partial_{t}$. Therefore, a photon space is a photon surface with constant $x(t)$, which in view of Proposition 2 is implicitly defined by those $x$ satisfying
@@
3 Spherical dust collapsing spacetimes
3.1 The global model
In the following, we will first review the basic properties of spherical dust spacetimes together with the features related to the nature of the spacetime singularities arising from complete collapse.
Let us consider a general spherical metric (2) where the energy–momentum tensor is given by $T=-\rho\,\mathrm{d}t\otimes\partial_{t}$ [Singh:1994tb, Joshi:2008zz]. Using Einstein field equations, the metric reduces to the Lemaître–Tolman–Bondi (LTB) model
@@ 3.1 The global model
The initial data $k(x)$ and $m(x)$ can be prescribed at some initial time $t_{0}$, i.e., $t_{0}=0$, along with the initial energy $\rho_{0}(x):=\rho(t_{0},x)$. From (8) the identity $m(x)=\frac{r}{2}\left(1-g^{\alpha\beta}r_{\alpha}r_{\beta}\right)$ is immediately derived: the function $m(x)$ is the so–called Misner-Sharp mass, and its choice will also determine the initial condition $r(t_{0},x)$, because integrating the Einstein field equation $m_{x}=4\pi\rho r^{2}r_{x}$ along $t=t_{0}$ one obtains
@@
The metric (7) describes only the interior region of the model, i.e., the dust star. In other words, this metric is defined for $x\in[0,x_{b}]$, where $x_{b}$ represents the boundary of the star. The timelike hypersurface $\Sigma=\{x=x_{b}\}$ must be matched with a suitable exterior solution that describes the matter distribution outside the spherical dust cloud. This will be achieved using the Darmois junction conditions, stating that the matching of two spacetimes across a timelike hypersurface $\Sigma$ is smooth if the first and the second fundamental forms of the two metrics in $\Sigma$ respectively coincide [Darmois1927, Israel:1966rt, Mars:1993mj]. It is well known that for the exterior region to be Schwarzschild spacetime (i.e., spherical vacuum), the radial pressure, indicated by the $T^{1}_{1}$ component of the energy-momentum tensor, must be zero at $\Sigma$. Naturally, this applies to the model being analyzed here, where pressures uniformly vanish throughout the spacetime. According to the Einstein field equations, this condition ensures that the Misner–Sharp mass remains constant across the matching hypersurface, a necessary condition for matching [Magli_1997].
It is worth observing that the construction of the global model—an LTB interior matched to a Schwarzschild exterior across the timelike hypersurface $\Sigma$—has the structure of an initial-boundary value problem (see e.g. [Brito:2017ijp]): the Darmois conditions play the role of boundary conditions that must be satisfied throughout the evolution, not only on the initial slice. In the present case, these conditions reduce to the requirement that the mass is continuous along $\Sigma$, together with the identification of the areal radius on $\Sigma$, and they are automatically propagated by the explicit solution of (8).
3.2 Final fate of the marginally bound dust collapse
@@
$\{(t,x)\,:\,x\in[0,x_{b}],\,\,t<t_{s}(x)\},$
matched to a Schwarzschild exterior
at the star boundary $\Sigma=\{x=x_{b}\}$. As discussed in the previous subsection, the Darmois junction conditions ensure that the matching is $C^{1}$ across $\Sigma$, provided that the Schwarzschild mass is given by $M=m(x_{b})$. Observe that the coordinates $(\tau,r)$ in (10) will be related to $t$ on the matching surface – in particular, $r=r(t,x_{b})$ on $\Sigma$ [^2] That justifies the slight abuse of notation where we employ the same letter $r$ for both the radial coordinate in (10) and the metric coefficient (9) within (7).. Details can be found in [Giambo2005GravitationalFields].
We will refer to this model as the marginally bound spherical dust collapse. Its end state, starting from regular initial conditions at $t=0$, is well known in the literature [Singh:1994tb]: let us summarize it briefly. Consider a mass profile such that, using (9)
@@
4 Photon surfaces in dust collapse
4.1 Extension to the interior LTB metric
Let us now determine the photon surface for the marginally bound dust collapse. On the exterior (10), of course, the photon surface is given by $r=3M$, see Example 1.
For the LTB interior metric, as pointed out in [Cao:2019vlu, (5.5)], a lengthy but straightforward computation transforms the equation (3) for photon surfaces as
that applying Proposition 2 can be given in the form of the following 1st-order system:
$\dot{x}(t)$ $=\frac{\sqrt{k(x)+1}}{r_{x}}\,y(t),$ (13a)
The crucial step now is to set initial conditions for the system (13a)–(13b). The first condition is necessarily given by $x(t_{0})=x_{b}$, where $t_{0}$ satisfies the equation $r(t_{0},x_{b})=3m(x_{b})$, i.e., using (9),
@@ 4.1 Extension to the interior LTB metric
Of course, one may consider other extensions that start spacelike at $x=x_{b}$. These choices correspond to setting $|y_{0}|<1$. However, $y_{0}<0$ would correspond to a decreasing $x(t)$ curve and would, for the same reason, be less significant than the $y_{0}\geq 0$ extension. The following lemma will help us rule out all the other cases.
@@
The marginally bound spherical dust collapse admits an extension of the photon surface $r=3M$ in the interior of the cloud. This is the only physically acceptable extension, and it is given by the surface
@@
4.2 Photon surface analysis
@@
The curve $t_{\xi}(x)$ is a supersolution of (18) and $x_{b}$ is chosen such that $t_{0}\in]t_{\xi}(x_{b}),t_{h}(x_{b})[$. The solution of (18) can be extended back for $x<x_{b}$: it cannot cross the subsolution $t_{h}(x)$ from below nor the supersolution $t_{\xi}(x)$ from above and then it must extend up to the central singularity $(x=0,t=1)$.
@@
In the marginally bound spherical dust collapse, the photon surface extends back to the central singularity if and only if it is naked. In this case, there exist null radial geodesics $t_{g}(x)$ emanating from the central naked singularity such that $t_{g}(x)<t(x)$, $\forall x\in]0,x_{b}]$, where $t(x)$ is the solution of (18) associated with the photon surface extension established in Theorem 3.
@@ 5 Discussion and conclusions
Our results indicate that the only physically consistent extension of the photon surface $r=3M$ into the collapsing cloud is given by a surface generated by outgoing null radial geodesics. This extension satisfies the photon surface equation and adheres to the junction conditions at the boundary with the Schwarzschild exterior. That is a static geometry, and this is precisely why the photon sphere $r=3M$ is timelike: the light cones at fixed r do not change with time, so circular null orbits can remain tangent to the timelike worldtube $r=3M$. Inside the LTB cloud, the metric coefficients depend on comoving time, and the light cones tilt progressively inward as the dust collapses. A timelike hypersurface playing the role of a photon surface would need to track these changing light cones without letting tangent photons shear away from it, which becomes impossible once the infall is strong enough (see Lemma 1). The photon surface, therefore, has to become null: in the time-dependent collapse here considered, the only way to preserve the trapping property is to let the surface coincide with a congruence of radial null geodesics.
Importantly, the nature of the central singularity—whether it is covered or naked—plays a crucial role in determining the global structure of the photon surface. When the singularity is covered, the photon surface terminates at the regular center before the singularity forms. Conversely, if the singularity is naked, the photon surface can extend all the way back to the central singularity itself. However, in such cases, the surface is not sufficient to prevent the escape of photons from the singularity, highlighting its limited covering capability. Previous studies have established that naked singularities create different lensing effects compared to black holes [Virbhadra:2022iiy]. Additionally, research by Ortiz et al. [Ortiz:2015rma] demonstrates differences in the shadow evolution between naked singularities and black holes.
While both objects eventually appear similar to distant observers, their formation dynamics differ significantly, also due to the infinitely redshifted photons that escape from the central naked singularity to the external region – a phenomenon documented in earlier work [Dwivedi:1998ts, Giambo:2006vr]. Our findings in the present paper support these observations by showing that shadow evolution also varies in terms of photon surface extension, suggesting that understanding photon surface geometry is crucial for distinguishing these compact objects, as the formation and features of the shadow are determined by the whole underlying spacetime structure.
From an observational standpoint, the key implication of Theorem 4 is that the early-time evolution of the shadow differs between the two end-states. When the central singularity is naked, the photon surface extends to the singularity itself but fails to prevent the escape of null geodesics (those satisfying $t_{g}(x)<t(x)$); the shadow, therefore, forms with a delay and grows more slowly than in the covered case, where the photon surface reaches the regular centre and the shadow develops monotonically. In principle, a time-resolved sequence of shadow images—or, equivalently, the spectral evolution of photons traversing the collapsing cloud—could capture this distinction. However, it must be remembered that the collapse timescale $\sim GM/c^{3}$, as one can infer from standard dimensional analysis – this fact and the opacity of realistic matter severely limit current observational prospects, making such tests a target for next-generation instruments rather than present facilities.
@@ Acknowledgement.
The authors acknowledge the support of INdAM. In particular, RG is partially supported by ”INdAM–GNAMPA Project” CUP E53C25002010001. The authors thank LM Cao and Y Song for their insightful discussions, and the anonymous Referees for their helpful remarks and suggestions.